OpenStudy (anonymous):

What value of k makes the equation true? 12 + (–17) = –17 + k k =

OpenStudy (anonymous):

12 because of commutative property of addition.

OpenStudy (f_jayyy):

12=-17+17+k 12=k k=-12 (negative 17 canceled out the positive 17)

OpenStudy (anonymous):

so?

OpenStudy (f_jayyy):

i just worked out the whole equation for you

OpenStudy (anonymous):

oh i get it now ok

OpenStudy (anonymous):

Think about it like this, Both sides have to be equal right? There's 1 number that is the same on both sides. There is another 2 numbers on one side and on the other a variable Well, if A+B=C Then, B+__=C ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the answer is 12